40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 143 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  9885   Wed Apr 30 21:31:25 2014 ranaHowToIOOMystery Alignment again

Looks like there was some mysterious MC alignment shift around 5:30 PM today, but no elog.....?? Now things are drifting much more than this morning or yesterday. Who did what and why???

I think I'll blame Jamie since he just got back and did some computer fiber voodoo today.

http://www.lawsome.net/no-throwing-rotten-tomatoes-a-repealed-kentucky-law/

  9884   Wed Apr 30 21:16:42 2014 ranaUpdateLSCMC2 Strad

bettsreplica.jpg

I found the YARM LSC feedback going to MC2 and the MC2 violin mode (at 644.69 Hz) rung up. The existing notch was just a second order Twin-T style notch (so not a good idea) and also not turned on, since it was in the FM4 spot of LSC-MC2 and the vio triggers are ganged between all mirrors and don't touch FM4.

I copied the PRM vio bandstop into FM2 of this bank, deleted the old notch, and tuned the bandstop frequencies a little to get the violin peak into one of the zeros of the elliptic bandstop. Attached are the Y-arm / MCF spectrum with the mode rung up as well as the new filter's TF compared with the old notch.

P.S. I installed http://en.wikipedia.org/wiki/Midnight_Commander on pianosa.

Attachment 2: MC_Y_vio.pdf
MC_Y_vio.pdf
Attachment 3: MC2_vio.pdf
MC2_vio.pdf
  9883   Wed Apr 30 18:06:06 2014 jamieUpdateCDSPOP QPD signals now on dolphin

The POP QPD X/Y/SUM signals, which are acquired in c1ioo, are now being broadcast over dolphin.  c1ass was modified to pick them up there as well:

c1ioo-POPQPD.pngc1ass-POPQPD.png

Here are the new IPC entries:

controls@fb ~ 0$ egrep -A5 'C1:IOO-POP' /opt/rtcds/caltech/c1/chans/ipc/C1.ipc
[C1:IOO-POP_QPD_SUM]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=116
desc=Automatically generated by feCodeGen.pl on 2014_Apr_30_17:33:22
--
[C1:IOO-POP_QPD_X]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=117
desc=Automatically generated by feCodeGen.pl on 2014_Apr_30_17:33:22
--
[C1:IOO-POP_QPD_Y]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=118
desc=Automatically generated by feCodeGen.pl on 2014_Apr_30_17:33:22
controls@fb ~ 0$ 

Both c1ioo and c1ass were rebuild/install/restarted, and everything came up fine.

The corresponding cruft was removed from c1rfm, which was also rebuild/installed/restarted.

  9882   Wed Apr 30 17:45:34 2014 jamieUpdateCDSc1ioo now on Dolphin network

For reference, here are the new IPC entries that were made for the ALS X/Y phase between c1als and c1lsc:

controls@fb ~ 0$ egrep -A5 'C1:ALS-(X|Y)_PHASE' /opt/rtcds/caltech/c1/chans/ipc/C1.ipc
[C1:ALS-Y_PHASE]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=114
desc=Automatically generated by feCodeGen.pl on 2014_Apr_17_14:27:41
--
[C1:ALS-X_PHASE]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=115
desc=Automatically generated by feCodeGen.pl on 2014_Apr_17_14:28:53
controls@fb ~ 0$ 

After all this IPC cleanup is done we should go through and clean out all the defunct entries from the C1.ipc file.

  9881   Wed Apr 30 17:07:19 2014 jamieUpdateCDSc1ioo now on Dolphin network

The c1ioo host is now fully on the dolphin network!

After the mx stream issue from two weeks ago was resolved and determined to not be due to the introduction of dolphin on c1ioo, I went ahead and re-installed the dolphin host adapter card on c1ioo.  The Dolphin network configurations changes I made during the first attempt (see previous log in thread) were still in place.  Once I rebooted the c1ioo machine, everything came up fine:

dolphin.png

We then tested the interface by making a cdsIPCx-PCIE connection between the c1ioo/c1als model and the c1lsc/c1lsc model for the ALS-X beat note fine phase signal.  We then locked both ALS X and Y, and compared the signals against the existing ALS-Y beat note phase connection that passes through c1sus/c1rfm via an RFM IPC:

The signal is perfectly coherent and we've gained ~25 degrees of phase at 1kHz.  EricQ calculates that the delay for this signal has changed from:

ALSXonDolphin.pdf

122 us -> 61 us 

I then went ahead and made the needed modifications for ALS-Y as well, and removed ALS->LSC stuff in the c1rfm model.

Next up: move the RFM card from the c1sus machine to the c1lsc machine, and eliminate c1sus/c1rfm model entirely.

  9880   Wed Apr 30 16:07:59 2014 manasaUpdateLSCALS X noise post servo modification

I. The Y arm stayed stable through last night and I have saved the arm lock settings to IFOconfigure.

II. ALS X arm noise measurements

I looked at the before and after noise of ALSX.

Settings:
Phase tracker gain = 85
Xarm servo gain = -17

The rms in loop noise has dropped from 3KHz to 500 Hz.

Attachment 1: Phase tracker OLTF

Attachment 2: Free running noise and in loop noise

Attachment 3: Out of loop noise (measured with arms locked using PDH for IR)

Attachment 4: ALS arm servo OLTF

xml data files can be found in /users/manasa/data/140430/

Attachment 1: ALSX_PToltf.jpg
ALSX_PToltf.jpg
Attachment 2: ALSX_FreeInloop.jpg
ALSX_FreeInloop.jpg
Attachment 3: ALSX_ool.jpg
ALSX_ool.jpg
Attachment 4: ALSX_OLTF.jpg
ALSX_OLTF.jpg
  9879   Wed Apr 30 14:21:50 2014 manasaUpdateCDSfb restarted

c1sus and c1isey were not talking to fb. The usual mxstream restart did not help.

Restarted fb

>>ssh fb

>>telnet fb 8087
shutdown

All lights on the FE status screen are green now.

Note that Steve did an mxstreamrestart earlier today because the same machines c1sus and c1isey were not talking to fb.

  9878   Wed Apr 30 11:16:41 2014 SteveUpdateIOOthis typical morning

c1sus and c1iscey were reset. The PMC needed to be locked. MC locked instantly.

The FSS ion pump power supply was turned on.

 

 

 

 

Attachment 1: 40days.png
40days.png
  9877   Wed Apr 30 00:40:55 2014 manasaConfigurationLSCY arm IR lock troubleshooting

[Koji, Manasa]

The Y arm locks stably for IR PDH now.  

The reason for ETMY getting kicked during lock acquisition was finally found to be related to the limiter value set in the Y arm servo. We reduced the limiter value unintentionally and found that the lock acquisition stayed smooth. The limiter value was stepped in 1000s from 7000 and eventually found that the ETMY suspension was kicked when we try to acquire lock with the limiter value was set at 11000. 

 

The limiter for X arm at 11000 is not causing any trouble at the moment.

In the process, we did a bunch of things through the evening to troubleshoot IR locking of the Y arm.

Earlier today running the IFO configure script did not restore the arms to lock and both the ETMs needed to be aligned to lock the arms. The arms stayed locked for 15 minutes and the Y arm lost lock eventually leaving the ETMY in a misaligned state. 

The state of the Y arm was similar to what Jenne has explained in ELOG where the ETMY was kicked during lock acquisition and would move to a misaligned state.

To trouble shoot, there were several things that were done. A few of them might not have any direct correlation to the locking issue but could just be a coincidence.  

1.  The trigger time for the filters in the arm filter modules were set such that they switch on after the SUS violin filters. Arm FM trigger time = 3 s (previously set at 0.1s) and SUS violin trigger time = 1s. This reduced the number of lock loss events.

2. There was some drop in transmission when the bounce filter of Y arm (FM6) turned ON. This was fixed by changing its ramp time (initially set at 1s). The filter has been modified to turn on immediately upon arm lock acquisition before the other triggered filters in the filter module turn on.

3. The QPD and SUS signal cables running to the rack were checked to be intact. Koji found some of them to be loose. But this had no evident correlation with the arm locking problem.

4.The oplev and PD alignment was checked at the Y end. The high gain trans PD for Y arm was checked for good alignment by looking at TRY. It was found that the EXIT light at the Y end is injecting some noise to the transmission PD. 

5. The ETMY was given offsets in PIT, YAW and POS and the OSEM sensor values were checked to see if the suspension is behaving well. It was behaving well.

  9876   Tue Apr 29 16:42:29 2014 SteveUpdateVACTP2 drypump replaced

Quote:

 

 TP2's fore line - dry pump replaced at performance level 600 mTorr after 10,377 hrs of continuous operation.

Where are the foreline pressure gauges? These values are not on the vac.medm screen.

The new tip seal dry pump lowered the small turbo foreline pressure 10x

TP2fl after 2 day of pumping 65mTorr

 TP2 dry pump replaced at fore pump pressure 1 Torr,  TP2 50K_rpm 0.34A

 Top seal life 6,362 hrs

 New seal performance at 1 hr  36 mTorr, 

 Maglev at 560 Hz, cc1 6e-6 Torr

 

Attachment 1: dryforepumpreplaced.png
dryforepumpreplaced.png
  9875   Tue Apr 29 10:01:25 2014 KojiConfigurationGeneralnetgpibdata is working again now

I've moved the WB network analyzer to the OMC lab. The 40m network analyzer is not in service for the MC monitoring.
I setup the configuration so that the same command gives us the same spectrum measurement.

  9874   Tue Apr 29 01:10:16 2014 KojiConfigurationLSCNew ALS servo implemented for the X arm

New ALS servo performance

Attachment 1:

Comparison between the old (orange) and new (red). The new error signal (red) is suppressed like a white noise as expected.

Comparison between the out-of-loop evaluation (black) and the in-loop signal (red). Below 50Hz the out-of-loop is limited by the sensor-noise like something.
This out-of-loop stability was measured with the ALS stayed at the top of the resonance and calibrated the POX11 error signal.

Attachment 2:

New ALS servo with the LSC PDH signal. When the PDH signal is used for the control, FM4 is additionally used.
In this condition, the error signal was measured and calibrated into frequncy noise (Hz/sqrtHz).

By comparing the POX (with the new servo) and POY (with the old servo) signals, one can see that the new servo has better supression below 30Hz with almost no cost at ~100Hz.

Attachment 1: ALSX_SPE.pdf
ALSX_SPE.pdf
Attachment 2: ALSX_PDH_SPE.pdf
ALSX_PDH_SPE.pdf
  9873   Tue Apr 29 00:11:03 2014 steveUpdatesafetysafety audit 2014

Be aware that this may affect POP QPD and POP RF Thorlabs PD

Quote:

 Last long extension cord removed from 1Y1 to ITMX-ISCT

The AC power strip at ITMX-ISCT is coming from wall L#26

 

  9872   Mon Apr 28 23:05:03 2014 JenneUpdateLSCALS CARM and DARM settings

[Jenne, Koji]

The IFO is being uncooperative tonight, and I have an early morning meeting, so I'm calling it a night. 

Koji's filter module changes have been propagated from the Xarm to the Yarm, to CARM and to DARM.  (Actually, Q overwrote the changes to Xarm on Sunday accidentally, so first he reverted those for us, and then we propagated the changes). 

Today, with careful measuring, we find that for X and Y arms individually locked with the ALS, we want the gains to be +17 for the Yarm, and -17 for the Xarm (with the beatnote up-is-up convention).  This puts the UGFs at 150 Hz. 

We then switched over to CARM and DARM locking.  We guessed that the gains should be a factor of 2 lower since we're pushing on both ETMs for DARM, and the MC2 actuator is roughly the same strength as the sum of the ETMs.  In the end, after measuring the CARM and DARM loops, we find that the gains should be +7.5 for CARM, and +8.0 for DARM to set the UGFs at 150 Hz.  The servo is a little bit delicate, so having too low of gain is not okay. 

For some reason, we seem to be utilizing more actuator range with the new setup, so the limiters in the filter banks have been set to 11,000 (previously were 8,000), and the ALS watch script (ALSdown.py) threshold has been increased to 10,000 (previously 7,000). 

When finding the IR resonances with the new scheme, we are having trouble holding lock throughout the scan.  I have set the tramp for the coarse part of the scan to be 0.05 seconds (previously 0.01 seconds), which is an increase of a factor of 5 in the ramp time.  This helps, but may still not be enough, since we don't always hold lock until both IR resonances are found.

Probably the most annoying thing from tonight is the fact that ETMY keeps drifting off, particularly in yaw, when locked.  I don't have an explanation of why this is happening, but you can watch it happen sometimes, and the lock will be lost shortly thereafter.  Definitely when we lose lock and the ETM gets kicked, it is far enough away in yaw alignment that I have to completely redo the Yarm alignment.  This happens whether or not the ETMY oplevs are on.

To summarize, 3 scripts have been modified:

(1) ALSdown - threshold increased  (Modification from last week - turns off the slow temp servos for the end lasers, clears histories)

(2) ALSfindIRresonance - increase ramp time

(3) Lock_ALS_CARMandDARM - final gain values set to 7.5 for CARM and 8 for DARM, no filters come on until gains all the way up, turns on new set of Koji filters. (Modification from last week - turns on the slow temperature servos for the end lasers)

  9871   Mon Apr 28 20:31:38 2014 ranaSummaryIOOMC2_TRANS QPD Servo trend

This is a 4-day trend. I don't see any difference here which is significant. My guess is that the MC_TRANS servo gain is so low that its not really doing anything.

I'll turn it on periodically this week and then on Monday people can look at the trend again to see if they can identify when the servo is ON and when its OFF.

Attachment 1: Untitled.png
Untitled.png
  9870   Mon Apr 28 17:27:29 2014 steveUpdatesafetysafety audit 2014

Quote:

Quote:

 

 We had our annual safety inspection today.  Our SOPs are outdated. The full list of needed correction will be posted tomorrow.

 

The most useful found was that the ITMX-ISCT ac power is coming  from 1Y1 rack. This should actually go to 1Y2 LSC rack ?

 Please test this so we do not create more ground loops.

 Annual crane inspection is scheduled for 8-11am Monday, March 17, 2014

 

The control room Smart UPS has two red extension cords that has to be removed: Nodus and Linux1

 Last long extension cord removed from 1Y1 to ITMX-ISCT

The AC power strip at ITMX-ISCT is coming from wall L#26

  9869   Mon Apr 28 15:47:57 2014 manasaSummaryIOOMC2_TRANS QPD Servo trend

Quote:

We turned on the MC2_TRANS paths for both PIT/YAW tonight.

I turned off the BLP200 and turned on the RLP7 (RLP always are better than BLP). G_PIT = -0.111, G_YAW = 0.111. On Monday, let's let Steve look at the trends and determine if this centering servo is bad or good.

The MC was showing slow but periodic alignment drifts and eventually unlocked around noon. I looked up the alignment trend (Attach: 2 day trend)

MC_TRANS_PIT_ERR and MC_TRANS_Y_ERR show that the MC_TRANS servo slowly drifted the IMC alignment causing it to lose lock from time to time (mostly in yaw).

To confirm that the drift was NOT due to off-centering in the MC2_TRANS QPD, I turned off the WFS servo, moved MC2 to recenter the trans beam on the QPD, and re-enabled WFS servo.

MC_TRANS path in WFS is still left enabled.

Attachment 1: MC_TRANSservoApr28.png
MC_TRANSservoApr28.png
  9868   Mon Apr 28 13:18:18 2014 JenneUpdateLSCLSC offsets script modified

Quote:

The weird jumps at the beginning of each TRX peak are due to the triggered switching between the Thorlabs trans PD and the QPD trans PD.  Clearly we need to work on their relative normalizations.  There are also little jumps after each peak as the triggering sends the signal back to the Thorlabs PD.

 I was unhappy with the discontinuities between the Thorlabs and QPD versions of our transmitted light powers.  I realized that in the olden days, we just used the Thorlabs PD, and we set the no-light offset in the LSC version of the TR[x,y] filter banks.  However, now that we have brought the QPDs back, we are setting the dark offsets in the end suspension models, so that the signal chosen by the trigger already has its offset taken care of before we send it to the LSC model. 

Anyhow, having the offsets script try to put a value in the C1:LSC-TR[x,y]_OFFSET was giving us an extra offset and then when we did the normalizations, the numbers came out all wrong.  So.  I have removed the C1:LSC-TR[x,y] filter banks from the offset list, since they were made redundant. 

I have redone the normalizations for both arms (after running the ASS scripts).  I checked by watching the _OUT16 versions of the Thorlabs and QPD diodes before the triggering happens, and as I put offsets into the LSC servos to change the transmitted power, the diodes both change in the same way.  So, we'll have to see if this holds true for more than just values 0-1 the next time we lock, but hopefully it won't need changing for a while.

  9867   Mon Apr 28 11:08:11 2014 KojiUpdateLSCNew ALS servo design: expected error signals

Here are the MATLAB scripts and LISO codes for all of these servo analyses

Attachment 1: 140421_ALS_servo.zip
  9866   Mon Apr 28 11:03:57 2014 KojiConfigurationLSCNew ALS servo implemented for the X arm

The new ALS/LSC servo was implemented for the X arm.

I'll upload more data later but here I make quick remarks:

- We need to give the gain of 12 to have correct UGF with the ALS.

- With this servo, the Xarm PDH lock oscillates with the gain of 0.02. We need to lower the gain to 0.015.
  Also FM trigger should be changed not to trigger unused FMs (FM7/8)

  9865   Mon Apr 28 10:59:54 2014 KojiUpdateLSCNew ALS servo design: expected error signals

The expected error signals derived from the estimated free running error signals of the ALS.

1) Previously estimated free-running noise (blue)
2) Previous in-loop ALS error signal (red)
3) Estimated error signal with the new servo (green)
4) Out-of-loop noise of the ALS with the arm controlled with the IR PDH (black)

Now the error signal (green) is expected to be very white.
The suppressed noise between 3 to 20Hz are below the sensing noise level.
There seems a little excess at 24.5Hz and 28Hz. If it is limiting the RMS, we need to take care of them.

Attachment 1: ALSX_SPE_new.pdf
ALSX_SPE_new.pdf
Attachment 2: ALSY_SPE_new.pdf
ALSY_SPE_new.pdf
  9864   Mon Apr 28 10:48:48 2014 KojiUpdateLSCnew ALS servo design: comparison

Comparison of the new and old servo OLTF
The new servo has the same UGF, slightly less phase margin, and more gain between 1.5 and 25Hz.

Attachment 1: ALSX_OLTF_new.pdf
ALSX_OLTF_new.pdf
Attachment 2: ALSY_OLTF_new.pdf
ALSY_OLTF_new.pdf
  9863   Mon Apr 28 10:34:51 2014 KojiUpdateLSCnew ALS servo design

Based on the evaluation of the error signals, the new servo was designed.

Concept:
- Don't touch the locking filters. (i.e. FM5)
- Sacrifice some phase at 150Hz to increase the gain between 3-20Hz.
- As resonant gains costs the phase without increasing the LF gains, replace them with a poles for the integrators.


FM(:,1) = zero2(f,.5,.7).*pole2(f,0.001,.7)*(0.5/0.001)^2;
FM(:,2) = zero2(f,5,2).*pole2(f,3,3).*pole1(f,1).*zero1(f,5)*5*(5/3)^2;
FM(:,3) = zero2(f,25,.7).*pole2(f,3.2,10)*(25/3.2)^2; % Zero crossing
FM(:,4) = zero2(f,35,2).*pole2(f,3,3).*zero1(f,3000).*pole1(f,1).*pole2(f,3000,1/sqrt(2)).*pole1(f,700).*zero1(f,10).*zero1(f,350).*136e1;
FM(:,5) = zero1(f,1).*pole1(f,4010).*pole2(f,17.3211e3,1.242).*zero2(f,18.865e3,100e3);
FM(:,6) = zero2(f,5,2).*pole2(f,10,2).*pole2(f,16.5,30).*zero2(f,30,2);
FM(:,7) = 1;
FM(:,8) = 1;
FM(:,9) = 1;
FM(:,10) = 1;
dc_gain = 14;

FM1/2/3/5/6 are expected to be used for the control.


FM1: Boost below 0.5Hz. This does not cost the phase margin.
FM2: Increase the gain below 5Hz. This hardly cost the phase margin.
FM3: Boost below 25Hz. This is the main phase cost at UGF. This has a complex pole pair at 3Hz with Q=10 to supress the stack motion.
FM6: zero-pole-pole-zero combination to boost the gain between 5 to 30Hz. This eats the phase margin a little.

Note that the phase tracker gain for the X arm was increased by factor of 2.5.

Attachment 1: ALSX_OLTF_new2.pdf
ALSX_OLTF_new2.pdf
Attachment 2: ALSY_OLTF_new2.pdf
ALSY_OLTF_new2.pdf
  9862   Mon Apr 28 10:24:10 2014 KojiUpdateLSCerror signal characterization

As we now have the loop model, we can characterize the error signals.

We have the following data:

1) Free-running ALS error signals (i.e. phase tracker output) calibrated in Hz (for 532nm) (blue)
2) Controlled ALS error signals calibrated in Hz (for 532nm) (red)
3) ALS error signals measured with X and Y arm locked with the IR PDH. (black)
    This is likely to represent the sensing noise of the beatnote detection

from 2) we can derive the similar quantity as 1)
4) Estimated free-running ALS error signals from the controlled signals (green)

Remarks:

- From 1) and 4) we can see that the phase tracker is not perfectly linear. It seems that fast fringing of the phase tracker is causing upconversion.

- From 2) and 3) the servo loops don't have enough gain between 3Hz and 20Hz. On the other hand they have too much gain bekow 3Hz.

Attachment 1: ALSX_SPE.pdf
ALSX_SPE.pdf
Attachment 2: ALSY_SPE.pdf
ALSY_SPE.pdf
  9861   Sun Apr 27 21:30:59 2014 KojiUpdateLSCALS servo characterization

The measured openloop TF of the ALS servo for each was characterized by a ZPK model.

The openloop TF can be modeled by:

1) Filter TF obtained from foton
2) Actuator response with appropriate assumption
3) Phase tracker closed loop TF
4) Delay caused by the digital control
5) anything else

For 1) ZPK models of the servo filter was obtained from foton. It turned out that the TF of FM5 doesn't match with the ZPK model in foton.
Therefore the TF was exported and fitted with LISO. This seems to be related to the pole frequency (3kHz) which is too close to Nyquist frequency (8kHz).

FM(:,1)  = zero1(f,5).*pole1(f,0.001)*5000;
FM(:,2)  = zero1(f,1).*pole1(f,0.001)*1000;
FM(:,3)  = zero2(f,4.5,1.4619).*pole1(f,0.001).*pole1(f,0.001)*20.2501*1e6;
FM(:,4)  = zero2(f,35,2).*pole2(f,3,3).*zero1(f,3000).*pole1(f,1).*pole2(f,3000,1/sqrt(2)).*pole1(f,700).*zero1(f,10).*zero1(f,350).*136e1;
FM(:,5)  = zero1(f,1).*pole1(f,4.010e3).*pole2(f,17.3211e3,1.242).*zero2(f,18.865e3,100e3);
FM(:,6)  = zero2(f,3.2,0.966775).*pole2(f,3.2,30.572);
FM(:,7)  = zero2(f,16.5,2.48494).*pole2(f,16.5,78.5807).*zero2(f,24.0,2.22483).*pole2(f,24.0,7.03551);
FM(:,8)  = 1;
FM(:,9)  = zero2(f,7.50359,1.07194).*pole2(f,1.43429,0.717146)*27.5653;
FM(:,10) = 1;

dc_gain = 14;

FM1/2/3/5/6/7/9 are used for the control.

For 2), a resonant freq of 0.97 with Q of 5 was assumed.

The model for 3) was obtained by the previous entry.

Now the measured TF was divided by the known part of the model 1) ~ 3) and empirically fitted in LISO.

### XARM ###
pole 392.5021429051 698.1992431753m
zero 42.3128869460k 31.0954443799m
pole 589.2716424428 2.8325268375
factor 8.3430140244
delay 34.7536691023p

### YARM ###
pole 416.2463334253 743.2196174175m
zero 97.9161062704M 114.6703921876m
pole 626.0463515310 2.7671041771

factor 9.0045911761
delay 34.0945727358p

These compensation TF have weird TF. Probably the frequency response of the delay and the analog AA/AI filters without the high frequency data
led the LISO make up this. I'm requesting Masayuki to provide the AA/AI data to make the estimation more reasonable.
For the servo modeling, this is sufficient and we'll go a head.

The results of the OLTF modeling are attached.

Attachment 1: ALSX_OLTF.pdf
ALSX_OLTF.pdf
Attachment 2: ALSY_OLTF.pdf
ALSY_OLTF.pdf
  9860   Sun Apr 27 20:26:19 2014 KojiUpdateLSCPhase Tracker servo characterization

The measured open loop TF of the ALS Phase Tracker loop for each arm was characterized by an empirical model on LISO.

The model for the open loop TF has pole 1m instead of the one at DC as LISO has a difficulty to model it.
Digital time delay and the sampling effect seem to be well represented by a zero at ~8kHz and delay of  ~60us.
(cf 16kHz sampling => 61us)

The XARM phase tracker has the UGF of 1.5kHz. This is too low because
1) The phase rotation at 100Hz is visible in the plot.
2) We don't much care about the closed loop bump in the phase tracker as long as the phase tracker keeps its continuity.

So I suggest to increase the gain so that we have the UGF of 3kHz. (phase margin: 24deg)

The red curve in the plot is the closed loop response calculated by CLTF =  - OLTF / (1-OLTF).

The model results are used in the ALS servo models.

Attachment 1: ALSX_PTTF.pdf
ALSX_PTTF.pdf
Attachment 2: ALSY_PTTF.pdf
ALSY_PTTF.pdf
  9859   Sun Apr 27 19:53:54 2014 ericqUpdateLSCPRFP YArm Locking

Inspired by a comment by Koji the other day, I spent some time yesterday and today working on locking a (very lossy) power recycled Y-arm. ITMX was misaligned, to save myself the headache of dealing with ITMY getting a sign flip and ITMX staying the same when the arm resonates. 

My main goal was to achieve high bandwidth control with the analog CARM servo. 

TL,DR: Transisitoned 90% to REFLDC through CM_SLOW at TRY = 2.1 twice. Couldn't make it all the way over. 


PRCL settings:

  • Input: REFL165 I.
  • Actuate on PRM +1
  • Control: G=-.32 (~100Hz UGF); Acq on FM 4,5; Trig 1,2,3,6,9 (I modified the +10dB in FM1 to a 1kHz ELP)
  • Trig: POP 110 I: 1.5 up, 0 down (max was around 4 counts, very weak PRC!)

The PRC was very stable in this configuration, which doesn't surprise me due to its simplicity. I was honestly a little surprised there was enough light to lock on 3f. REFL33 didn't work. 


My efforts to bring the Y-arm into lock were very similar to the CARM procedure we've been using recently. (Which is the motivation for this exercise)

At first I was actuating on ETMY, and got to the point where I wanted to start bringing in the CARM servo slow output, then realized that I didn't want to actuate both on the ETM and MC AO. (Maybe this would be doable, but in the end, not what I'm interested in learning about in terms of overlap with CARM locking)

From then on, I only actuated YARM on MC2. (Heads up, my lock-losses will show up in the trends of the MC2 Trans addition to the WFS.)

Transitioning the arm to SqrtInv TRY control was just as straightforward as it has been for CARM. However, engaging the LSCLock FM (FM4), would sometimes work beautifully, and sometimes kick the hell out of MC2. Keeping an eye on the error signal spectrum and UGF gave no indication which outcome would happen. Once FM4 could be engaged, the transmitted power was very stable. Without FM4, reducing the offset didn't get very far without losing lock. 

I tried a few times to bring in CM_slow (set to just IN2, i.e. offset adjusted REFLDC), at arbitrary arm powers, with little success. I didn't know how much arm power to expect at resonance, and thus didn't really know where on the line width I was.

I knew I was mostly outside of the linear regime of the PDH signals, since, even though I had good coherence between, say, REFL11 I and SqrtInvTry, with an ETMY excitation on; when I would turn TRY normalization on/off, I would see the sign of the TF change. 

I then realized that I could actively keep an eye on the trend of POY11, to see when I got to the PDH "hump", which is where REFLDC starts being usable, and SqrtInv is reaching its limit. 

This brought me to a YARM offset of .115, with a steady TRY of about 2.1. I adjusted the analog offset of the REFLDC input to the CARM board, and the digital gain of the CMSLOW input filter to get 1:1 correspondence between CMSLOW and the SQRTINVY channels. Their spectra were neigh identical, with CMSLOW having slightly more high frequency noise. 

I started stepping SQRTINV down by .1, and upping CMSLOW by .1. This shifted the offset around, so I opted for taking away gain before bringing it back, because I didn't want to get so close to resonance that SQRTINV would freak out. I got to .1*SQRTINVY + .9*CMSLOW, and lost lock. TRY was getting noisier as I made the transition. 

I'm not sure what exactly was the reason for failure. I'm going to go back over some of the data to try to get an idea.. Maybe I should've loosened up some of the gain/boosts during the transition. 


So, no great success story yet, but this configuration is a lot simpler than the full PRFPMI, and I feel that I should soon be able to get it fully controlled, and figure out a systematic way to make the digital to analog transition for this PRFP cavity, and thus have a much more informed basis for doing the same for CARM control. 

  9858   Sat Apr 26 13:19:59 2014 ericqSummaryIOOMC2_TRANS QPD Servo re-re-engaged again

Quote:

We turned on the MC2_TRANS paths for both PIT/YAW tonight.

I should've included this in my Thursday night ELOG... That evening, I aligned the mode cleaner with reasonable MC1/3 spot positions, and the MC2 spots very close to centered, and recentered the WFS and MC2 Trans QPDs. The mode cleaner held up very well over the course of that evening, even when actuating CARM on MC2 with WFS engaged (which previously wasn't very stable when the WFS weren't well aligned).

  9857   Fri Apr 25 23:08:57 2014 ranaSummaryIOOMC2_TRANS QPD Servo re-re-engaged again

We turned on the MC2_TRANS paths for both PIT/YAW tonight.

I turned off the BLP200 and turned on the RLP7 (RLP always are better than BLP). G_PIT = -0.111, G_YAW = 0.111. On Monday, let's let Steve look at the trends and determine if this centering servo is bad or good.

  9856   Fri Apr 25 22:20:01 2014 ranaUpdateIOOcsh/tcsh hackery combatted

To make the mcwfson/off scripts work from rossa (and not just Jamie's pet machine) I swapped the sh-bang line at the top of the script to use 'env bash' instead of 'env csh' in the case of mcwfsoff and 'env tcsh' in the case of mcwfson.

The script was failing to work due to $OSTYPE being defined for pianosa csh/tcsh, but not on rossa.

During debugging I also bypassed the ezcawrapper for ezcaswitch so that now when ezcaswitch is called, it directly runs the binary and not the script which calls the binary with numerous retries. In the future, all new scripts will be rewritten to use cdsutils, but until then beware of ezcaswitch failures.

WFS scripts checked into the SVN.

This was all in an effort to get Koji to allow me to upgrade pianosa to ubuntu 12 so that I can have ipython notebook on there.

Objections to upgrading pianosa? (chiara and megatron are already running ubuntu 12)

  9855   Fri Apr 25 13:18:08 2014 Dark JamieUpdateLSClocking activity

Quote:

[ericq, Jenne, Zach]

We spent some time tonight trying to push our CARM locking further, to little avail. DARM/CARM loop oscillations kept sneaking up on us. We measured some MC2 motion -> REFL11 Transfer Functions to see if we could see CARM plant features; plots will come in the near future...

 Probably things would have worked better if you would have gotten your hair done at the same place as me.

Attachment 1: m10008_f1_bg.jpg
m10008_f1_bg.jpg
  9854   Fri Apr 25 10:43:57 2014 KojiUpdateLSC(Fixed) Y end whitening board

I went to WB and found the last spare module of D990399 revB. We need to thank Frank for his foresight.

The original (=broken) board had various modifications from this revB.
I had to check the schemaric diagram and the difference between the boards and migrate some of the SMD components from left to right.


Here is the deciphered features of the QPD whitening board:
- The input stage is a VGA amp (AD602). It has the internal input impedance of 100 Ohm. The series resister
  of 909 Ohm gives us 1/10 voltage division! It is more tricky as the QPD (D990272) has the output impedances of 50Ohm
  (for the both side of the differential out) and on resistance of MAX333A. So it could have been deviated by ~10% from the nominal.

- Variable gain control: The input has 1/10 voltage division. The gain is fixed at the unity. In total the gain of the variable control stage is 1/10.
  This gives us the gain range of +42dB/-22dB for +10V/-10V. The actual range is limited to be -10~30dB.

- Whitening stages. Each channel has two sets of the whitening path and the bypass path.
  They could be switched by binary control inputs but I permanently enabled the whitening by pulling the MAX333 control inputs to the ground.
  The whitening zero and pole are at 4.02Hz and 40.6Hz.

  Each bypass path has an additional cap of 220pF in parallel to 35.7kOhm (R101 and R103 for CH1), resulting in the pole at 20.2kHz.
  Each whitening paths had a 5.6nF cap (C53 and C64). This cap was replaced with 350pF, resulting in the move of the pole freq from 800Hz to 12.7kHz.

- There are two anti-aliasing stages which were designed for 2kHz sampling rate. They are identical sallen key 2nd-order LPFs with fc=766Hz and Q=0.74 (~ butterworth).
  As all of these caps were removed, they are just voltage followers now.

- The final stage (AD620) has the gain resister of 16.5k. The gain is 1+(49.4k/16.5k) = 3.99.

- The 4pin lemo connector (J8) was removed from the board. We instead installed an isolated BNC connector on the panel for the thorlabs PD serving as the high gain PD.

- There is a daughter board for the high gain PD. This seems to be the butterworth low pass filter with fc=~30kHz.
  The differential output of the daughter board is connected to pin 17 and 18 of J10 (S5 Out and Rtn).

- The input of the daughter board is differential (AD620). Therefore the LEMO connectros next to the BNC were wrapped with Kapton tapes for isolation.

Board test at the workbench.

- The test required two dual power supply as the unit requires +/-5V and +/-15V.

- The four channels were tested with the signal injection. 1kHz input yielded 20mVpp across the AD602 input. The output of the 1st whitening stage was
  60mVpp. This makes sense as the gain of the AD620 is -10dB (1/10 and 10dB). The output of the 2nd whitening stage was 600mVpp.
  Finally the output of the output stage was confirmed to be 2400mVpp. This was confirmed for four channels.

- The daughter board output was also checked. The gain is the unity and flat upto ~10kHz.

Board installation

- Jenne installed the module. This time there was no smoke.


Gain mystery

- It was not sure how the whitening gains have been given.

- The corresponding database entry was found in /cvs/cds/caltech/target/c1auxey/ETMYaux.db as

grecord(ao,"C1:ASC-QPDY_S1WhiteGain")
grecord(ao,"C1:ASC-QPDY_S2WhiteGain")
grecord(ao,"C1:ASC-QPDY_S3WhiteGain")
grecord(ao,"C1:ASC-QPDY_S4WhiteGain")

- The gains for S2-S4 were set to be 30. However, C1:ASC-QPDY_S1WhiteGain was set to be 8.62068.
And it was not writable.

- After some investigation, it was found that the database was wrong. The DAC channel was changed from S100 to S0.
The corrected entry is shown here.

grecord(ao,"C1:ASC-QPDY_S1WhiteGain")
{
        field(DESC,"Whitening gain for QPDY Seg 1")
        field(DTYP,"VMIVME-4116")
        field(OUT,"#C0 S0 @")
        field(PREC,"1")
        field(EGUF,"42")
        field(EGUL,"-22")
        field(EGU,"dB")
        field(LINR,"LINEAR")
        field(DRVH,"30")
        field(DRVL,"-10")
        field(HOPR,"30")
        field(LOPR,"-10")
}

- Once c1auxey was rebooted, the S1 whitening gain became writable. Now all of the channels were set to be +30dB (max).

 

Attachment 1: D990399-B_40m.pdf
D990399-B_40m.pdf D990399-B_40m.pdf D990399-B_40m.pdf
Attachment 2: P4245552.JPG
P4245552.JPG
Attachment 3: P4245553.JPG
P4245553.JPG
Attachment 4: P4245551.JPG
P4245551.JPG
  9853   Fri Apr 25 03:14:46 2014 ericqUpdateLSClocking activity

[ericq, Jenne, Zach]

We spent some time tonight trying to push our CARM locking further, to little avail. DARM/CARM loop oscillations kept sneaking up on us. We measured some MC2 motion -> REFL11 Transfer Functions to see if we could see CARM plant features; plots will come in the near future...

  9852   Thu Apr 24 23:55:31 2014 KojiUpdateLSCY end whitening board

The main problem was a panel fixing bolt that caused the short circuits between power supply layers.
This burned the PCB and secondarily caused permanent short circuit between +15V/-15V/+5V layers.

Diagnosis

- The resistances between +15V, +5V, and -15V were low. The resistance between +15V and -15V is 13 Ohm.
  The one between +5V and -15V is 7Ohm. And the one between +15 and +5 is 19Ohm. So the situation is

                o -15V
                |
+15V o-(13 Ohm)-+-(9 Ohm)-o +5V

Even after removing all of the active components from the board, they remained the same.

- The tantalum caps were removed from the board and it was confirmed that they are not the cause of the issue.

- The panel was removed from the module for the component migration to a spare board (to be described in the other entry).
I found that the screw hole and the screw have burnt marks. The screw need an insulation tube to avoid short circuit.
The other screw was also bare. The spare board has the screws with the insulation tubes.

 

Attachment 1: P4245550.JPG
P4245550.JPG
  9851   Thu Apr 24 22:39:14 2014 ranaConfigurationGeneralnetgpibdata is working again now

Quote:

Not working again. I tried the commands in Koji's elog as well as the ones from my notes, but the AG 4395 hooked up to the yellow box named crocetta doesn't work. It gets to the stage of opening the data files and hangs. I tried it with many variations on pianosa and rossa. Also tried power cycling the analyzer and the Prologix and the bridge.

While hooked up to the MC error point I set the modulation frequency from 137 to 133 Hz to minimize the 3 MHz peak as usual.

Started working tonight. Don't know if anyone did anything to the Martian network or not, so its a mystery...

I also modified the script and SVN'd it. It now correctly takes your plot wants and adjust the linear/log of the axes accordingly

./SPAG4395A.py -i crocetta -A -v 4 --att=0 --start=1kHz --end=10MHz --bw=1kHz --plot --semiy

and also saves the plot as out.pdf

In the attached image I show the MC board TP1A output. The two peaks around 3.7 Mhz are the sideband beats we speak of. The lower one is proportional to the MC length/frequency mistmatch.

Attachment 1: out.pdf
out.pdf
  9850   Thu Apr 24 16:25:31 2014 ericqUpdateLSCQuick CM servo prep

I added ~1m of cable to the LO side of the REFL11 Demodulator, which brought its PRCL demod phase to about 8 degrees. According to my simulations, PRCL and CARM have the same angle (but opposite sign) at resonance. There seems to be a severe lack of SMA cables in the lab, so I didn't tune it to be any closer. Cos(8 degrees)=.99, so I think it should be fine to use it for the CARM servo, since none of the other signals are going to be nearly as big. I plugged analog REFL 11 I back into the CARM servo IN1. 

As for IN2, I threw together a temporary setup for using REFLDC as a complementary signal. I T'd off the REFLDC signal (which is the DC signal out of REFL55), and sent it into an SR560 to subtract an offset. The offset comes from a 1Hz-passive-pomona-box-low-passed C1:IOO-TT4_LR output, since there are 8 DAC channels set up for the nonexistent tip tilts 3 and 4 actively running. The output of the SR560 is sent to the CARM servo IN2. 

I adjusted the offset by turning on only IN2 in the CARM servo MEDM screen, and looking at the CM_SLOW signal in data viewer. I adjusted gains and such to get it to look just like REFLDC with the PRC locked. There was good coherence and no appreciable phase difference from DC out to some kHz, albeit a dip in coherence to about .8-.9 from ~40 to 300Hz, for some reason. (This included turning on the unWhite FM in the REFLDC filter bank)

If this signal turns out to be useful, it will be relatively straightforward to put together a little box that does the offset subtraction nicely, but this should do for our immediate needs.

Lastly, I hung up this plot in the control room to give us information about the DC values of different signals as the CARM offset changes. This is helpful to see what our CARM offset is, based on the transmission we se, when different signals start to have length dependence, where they start/stop being linear, etc. The TRX curve is scaled to a maximum of 600, REFLDC is normalized to input power = 1, and all the rest are arbitrarily scaled to fit on the plot. I've assumed 75ppm loss on all mirrors in my simulation (PRM, BS, 2xITM, 2xETM), mostly to get some realistic profile of REFLDC. 

carmOffsetProfiles.pdf

  9849   Thu Apr 24 14:23:09 2014 not manasaUpdateLSCY end whitening board

 

 maybe the tantalum caps on the daughter board power supply lines are blown? If so, replace with 35V+ ceramic.

  9848   Thu Apr 24 14:00:42 2014 JenneUpdateLSCLocking without TRY

Here is 1 second of data, with REFLDC, POPDC and TRX:

REFLDC_1sec.png

Here is a zoom of the first 3 big peaks in TRX.  The weird jumps at the beginning of each TRX peak are due to the triggered switching between the Thorlabs trans PD and the QPD trans PD.  Clearly we need to work on their relative normalizations.  There are also little jumps after each peak as the triggering sends the signal back to the Thorlabs PD.

REFLDC_3peaks.png

Here is a zoom of the single big peak about halfway through the 1 second of data:

REFLDC_1peak.png

And here is a zoom of the tail of that peak.  It looks to me like we want to start thinking about using REFL DC when our transmitted powers are around 2 counts.  We could do as soon as 1 count, but 2 is a little farther into the dip.

REFLDC_1peak_zoom.png

  9847   Thu Apr 24 11:19:50 2014 KojiUpdateLSCLocking without TRY

This seems the ever best stability at the zero offset PRFPMI.

Can you look at REFLDC in this data stream too? How was it promising?

  9846   Thu Apr 24 02:12:05 2014 JenneUpdateLSCLocking without TRY

I tried some locking anyway tonight, even though we don't have TRY. 

The biggest conclusion is that I miss the auto-resonance-finding.  I've been roughly scanning the Y-ALS offset to find the POY zero crossing when I see the resonance on the test mass face cameras. 

The next-biggest conclusion, is that I can hold the PRFPMI close to resonance, using ALS for CARM and DARM.  I was trying to transition DARM to AS55, but I couldn't get the last bit of the way.  That is, I couldn't turn off the ALS control.  So, I think that AS55 wasn't actually holding DARM, until maybe the last moment or so.

Anyhow, here are some time series.  My average TRX value is around 40 counts, and POPDC is maybe 250 counts (just PRMI, POPDC is about 75 counts).  Obviously this is noisy as hell, but I'm not using any IR signals for the arms.  Near the end of this first time series, I am trying to switch to AS55 for DARM.

TRX_avg_40cts_POPDC_avg_200cts.png

Zooming in, my real lockloss is due to PRCL oscillating at ~350 Hz:

Lockloss_PRCL_350Hz.png

However, I also saw ~25Hz peaks in CARM and DARM on the spectra starting to show up, and I see a ~25 Hz oscillation in DARM a few moments after the PRCL lockloss.  (Plot #2 is a zoom of the ~1.1 second mark on Plot #3.)

Lockloss_DARM_20Hz.png


The locking parameters:

CARM:

Input:  Using the new CESAR matrix, -1*ALSX, +1*ALSY.  Beatnotes both move up in freq if temp sliders move up.

Servo: gain = 6, FMs 1, 2, 3, 5, 6, 7, 9 on.  Offset = 0 counts. 

Output = -1*MC2

DARM:

Input:  +1*ALSX, +1*ALSY

Servo:  gain = 4.  FMs 1, 2, 3, 5, 6, 7, 9 on.  Offset = 0 counts.

Output = -1*ETMX, +1*ETMY

PRCL:

Input:  +1*REFL33_I, Norm = +0.01*POPDC, sqrt engaged.

Servo:  acquisition easier with -0.04 or -0.06, less gain peaking at -0.02   FMs 4, 5 on; 2, 3, 6, 9 triggered with 0.5 sec delay.  Servo trigger = POPDC, up 100, down 10.  FM trigger = POPDC, up 300, down 20.

Output = +1*PRM

PRCL ASC off, PRM oplev on.

MICH:

Input:  +1*REFL33_Q, Norm = +0.01*POPDC, sqrt engaged.

Servo:  gain = 2, FMs 4, 5 on; 2, 3 triggered with 0.2 sec delay.  Servo trigger = POPDC, up 100, down 10.  FM trigger = POPDC, up 300, down 20.

Output = +0.5*BS, -0.2625*PRM

PDs:

REFL33 analog gain set to 30 dB for both I&Q.

AS55 set to 0 dB for both I&Q.  AS55 had DC normalization of 80 counts (which was the measured number for PRFPMI when TRX was about 0.1 count this evening)

 

  9845   Thu Apr 24 00:11:35 2014 JenneUpdateLSCYend shutter back.

Quote:

To see if perhaps the shutter was the problem, I turned off the power to the Yend green shutter, and unplugged the cable.  The cable is laying on the table, with the connector sitting on a piece of plastic to isolate it.  Removing the shutter from the system did not change anything.

 I re-plugged in the Yend shutter, and turned it on.

  9844   Wed Apr 23 23:48:30 2014 manasaUpdateLSCY end whitening board

The MON outputs of the Y end QPD whitening board were hot earlier today while pulling it out of the crate. After swapping the 4 pin lemo connector with an isolated panel mount bnc connector, I stuck the board back into the crate and this immediately kicked the ETMY suspension. Jenne and I went to the Y end to look at what was going on. We removed the board from the crate after smelling something burning. The MON output ports of the whitening board were super hot this time. There is no sign of any components melting on the board (comparing the board with its pictures that were taken earlier) and a tester board stuck into the crate lights up just fine.

So the back panel is still ok. We need to troubleshoot or replace the whitening board.

Edit, JCD:  The attached photos are from right after I replaced the "Rgain" resistor, elog 9823.  What they show is that it looks like some of the melting / burning may have already been happening before I pulled the board, and I just never noticed :(  In particular, look at the resistors on the main board above the blue "G" sticker.  There isn't a difference that I can tell between this photo from last week, and today's situation. 

 

 IMG_1378.JPG

Attachment 1: IMG_1378.JPG
IMG_1378.JPG
Attachment 2: IMG_1379.JPG
IMG_1379.JPG
  9843   Wed Apr 23 19:58:00 2014 manasaUpdateLSCTRY 60Hz noise

[Steve, Manasa]

To find noise source

1. Swapped the power cable of the PD and checked that it is connected to the right power source.

2. Changed the aluminium base of the post holding the diode so that the diode is floating

3. Grounded the table and the rack

4. Routed the cable on the other side of the beam tube to isolate it from other cables.

After all the above, we still found that shaking the cable was making TRY noisy.

I pulled out the PD whitening board to replace the 4 pin lemo connector with a bnc connector so that we can swap the cable with a new one. So there is no TRY right now.

 

  9842   Tue Apr 22 22:49:10 2014 ranaUpdateLSCTRY 60Hz noise

 

 The detectors and electronics on this table are not properly isolated. To reduce the 60 Hz and ground loops, photodiodes and shutter must be isolated by using plastic spacers as we usually do elsewhere - this table just seems to have a few oversights.

Steve can start assembling all of the pieces to do this in the morning and then we can start the swapping after the meeting.

The high gain Transmon cable should be a regular BNC. There's no need for 4-pin LEMO in this usage, so the best move is to modify the board and replace the 4-pin LEMO connector with an isolated panel mount BNC female.

The AC adapter for this diode (and all of the detectors on the table) should get their power from a power strip which gets plugged into the rack with the whitening boards. The SHG oven, the Uniblitz shutter, and any cameras can get their power from another power strip if needed/wanted.

  9841   Tue Apr 22 21:54:50 2014 manasaUpdateLSCTRY 60Hz noise

Quote:

Quote:

 

P.S. I realigned the Y green to the arm and brought GTRY to 0.93

This evening, I was not able to successfully transition CARM from ALS to 1/sqrt(trans) signals.  The TRY time series looked odd, so I took a spectra, and we have huge 60Hz noise in TRY. 

Manasa, can you please take a look, and see if you can figure out what is going on?  We need TRY so that we can transition to 1/sqrt(trans) signals for CARM.  Thanks!!

 I went to the Y end to look at the TRY 60Hz noise situation this morning. While looking at TRY noise on dtt, I found that just lifting the cable away from the cable bunch that runs out of the table suppressed the noise drastically. 

Attachment 1

I removed the unwanted bnc connector in the path of the already long TRY cable running from the PD to the 1Y4 rack and isolated it from the bunch. TRY became less noisy.

But the noise was back again earlier in the evening and it looks like the noise is very much related to the TRY cable. TRY cable might have moved from its sweet spot while I was around checking cable connections yesterday.

I couldn't find a spare to replace it right away today (We need a BNC to 4 pin lemo).

Attachment 1: 60HzTRY.jpg
60HzTRY.jpg
  9840   Tue Apr 22 02:14:55 2014 ericqUpdateLSClock acquisition path for the CM servo

In an effort to familiarize myself with the analog CM servo, I've begun to replicate Koji and Den's work from the ELOG post that this is a reply to.

I hooked POX11Q into the IN1 of the CM board. (POX is rotated by ~86 degrees in the CDS, meaning analog Q is almost perfect.)

While there, I took out the too-long delay cables Jenne introduced for REFLs 11 and 55. (Also note: when we do cable-based analog phasing in the future, we should do it on the LO side, instead of the PD input side.) I also heard a dangerously crinkly sound from the short SMA cable for REFL11, so I replaced it with a beefy looking new one I found on the SP table.

I messed with the gain and offset in the CM_SLOW input filter to get it to look just like POX11_I_ERR, and was able to lock the arm on it without an issue. I then put the SR560 between the CM and MC (30k pole, but also AC coupled, because I figure the digital loop should be doing the work down there, and don't want to kick the AO with an offset), and was able to turn on the AO path with a gain of 8dB on the CM board and 10dB on the MC board, as detailed in Koji's procedures.

I wasn't able to increase the AO gain to 9dB without breaking lock, but maybe this is ok, because by judging by the LSC filter gains, POX11 might be about 3 times bigger than POY, so maybe 8dB AO gain on POX ~ =18dB AO gain on POY? I was able to put the CM servo offset at 0, but turning on boosts promptly kicked the MC out of lock.

I'm stopping for the night; but tomorrow I'll bust out a spectrum analyzer to see if I actually have won some bandwidth with the CM servo, and check out the situation with the offsets and boosts.

  9839   Tue Apr 22 01:39:57 2014 JenneUpdateCDSFB unhappy again

[Jenne, Q]

The frame builder (or something) is unhappy again.  I know that we've seen this before, but I can't find the elog entry that relates to this particular problem.

Every few minutes, the fb status lights on the CDS_STATUS screen go white, and then come back green.  It's annoying when it happens every hour or so (which is unfortunately typical), but it's pretty debilitating when it stops dataviewer and dtt every few minutes.  Just from the way the lights change, it looks like perhaps the daqd process is restarting itself periodically? 

  9838   Tue Apr 22 01:11:42 2014 JenneUpdateLSCTRY 60Hz noise

Quote:

 

P.S. I realigned the Y green to the arm and brought GTRY to 0.93

This evening, I was not able to successfully transition CARM from ALS to 1/sqrt(trans) signals.  The TRY time series looked odd, so I took a spectra, and we have huge 60Hz noise in TRY. 

I found a lock stretch from around 6:30pm that did not show the 60Hz noise, and then there was a lock stretch around 8pm that did have the noise.  So, something happened at the Yend between 6:30 and 8pm tonight.

Asking around, this was the time frame in which Manasa was down at the Yend to realign the green beam, and to check cabling for the PZT_OUT and ERR_MON signals to the ADC.

Looking at the spectra, Rana noted that we have even as well as odd harmonics of the 60Hz line, which is unusual.

TRY_60Hz_noise_21Apr2014.pdf

To try to diagnose the problem, Rana and I tried to make sure no cables' connectors were touching, and that no equipment was plugged in that shouldn't be.  We noticed that none of: the shutter, the Thorlabs TRY PD, or the QPD TRY are isolated from the table.  To see if perhaps the shutter was the problem, I turned off the power to the Yend green shutter, and unplugged the cable.  The cable is laying on the table, with the connector sitting on a piece of plastic to isolate it.  Removing the shutter from the system did not change anything. 

We don't see the 60Hz noise in the Xarm, so it's not on the laser light itself.  Also, we don't see the 60Hz lines in the Yarm feedback signal, so we're not putting the lines onto the mirror, and thus onto the Yarm's light. 

Manasa, can you please take a look, and see if you can figure out what is going on?  We need TRY so that we can transition to 1/sqrt(trans) signals for CARM.  Thanks!!

  9837   Mon Apr 21 23:33:57 2014 ranaSummaryGreen LockingHP 8591E reads low by 140 Hz out of 10 MHz

To check the basolute frequency stability of the old monochrome HP 8591E RF Spectrum analyzer that we're using for the ALS beat readout, I hooked its 10 MHz reference output (from its rear panel) into the A channel of the SRS SR620 frequency counter. The SR620 is locked to the FS 720 Rubidium clock via the 10 MHz connections in their rear panels.

So, we can assume that this is a good absolute readout. It reads 9.999860.7 +/- 0.3 Hz. So its 139.1-139.4 Hz lower than 10 MHz. The +/- 0.3 is just a slow drift that I see over the course of 10 minutes.

So, let's say that the analyzer is low by 10 ppm, so the arm length estimates are short by ~0.4 mm. A negligible correction, so there's no need to use atomic clocks to measure our arm lengths.

  9836   Mon Apr 21 22:53:16 2014 manasaUpdateLSCALS noise

Quote:

Last night, as well as tonight, the ALS seems not quite as robust as it was earlier in the week.

I have just taken noise spectra, and ALS is definitely more noisy than usual. 

These plots are with the arms held in CARM and DARM mode, with servo gains of 8. I was seeing the beginnings of gain peaking at a gain of 10, so I turned it back to 8.  Our ALS in-loop RMS is usually something like a few hundred Hz, but I'm seeing over 1kHz, so I have a factor of 4 or 5 too much noise.  Why?!?!?

I have noticed that ALS noise has been at 1KHz rms since LSC arm lock servos have been used to lock arms using ALS error signals. May be this has not been given much attention.

But looking more closely at the ALS noise (better dtt resolution for noise power spectrum) , there seems to be too much noise suppression at <1Hz and not much happening at around 10Hz.

Attachment 1 (data files at /users/manasa/data/140421/)

 

So I made a bunch of transfer function measurements for ALS and phase tracker servo. Koji will be using these and redesigning the servo filters so that we can get more suppression at 10Hz.

Other than this I also found that the Y arm showed more high frequency noise as compared to the X arm. (Edit by manasa: Thinking back now, this could be related to the onset of 60Hz noise at the Y end elog 9838. But still has to be looked at after fixing TRY)

Attachment 2

Tip: Once the arms are ALS locked, enabling the SLOW_SERVO helps hold the lock stably. 

P.S. I realigned the Y green to the arm and brought GTRY to 0.93

To do:

Find out what makes Y arm in-loop noise at high frequency higher than X arm.

Attachment 1: ALSX_FreeInLoop.jpg
ALSX_FreeInLoop.jpg
Attachment 2: ALSXY_inLoop.jpg
ALSXY_inLoop.jpg
ELOG V3.1.3-